# Distilling Multi-Level X-vector Knowledge for Small Footprint Speaker Verification

### 15 Dec 2023 v2

@@ IV-B Frame-level Embeddings
These types of embeddings share similarities with those from models such as d-vector [31]. The difference lies in the fact that, in this context, since the teacher model is an x-vector, during teacher training, utterance-level knowledge is back-propagated to the frame-level layers.
Additionally, for these frame-level embeddings, we explore the impact of mean normalization [55]. Specifically, we implement windowed mean normalization with a window size of $s=300$ frames and observe its normalization effect on the embeddings.
IV-C Aggregated Embeddings
Aiming to further exploit the potential of frame-level representations, drawing inspiration from [54], we adopt a strategy that involves aggregating the output from multiple frame-level layers via pooling methods to utterance-level. The resulting embedding is then obtained by taking the equal-weighted average of these aggregated outputs. Denoting $\boldsymbol{x}_{t}$ as the output teacher embedding, this approach is expressed as follows:
@@ IV-C Aggregated Embeddings
We benchmark the efficacy of two variants for $\mathbf{\Phi}(.)$. The first operator is statistics pooling (SP) [30], wherein the mean and standard deviation of $\boldsymbol{x_{i}}$ are computed and concatenated. The second operator is learnable dictionary encoding (LDE) [56], involving the aggregation of frame-level information from the TDNN layers through a convolutional and a linear encoding layer. The initialization of related weight values in LDE is accomplished using a normal distribution [57]. The corresponding embeddings derived from these operators are denoted as SP Aggr and LDE Aggr, respectively.
V Experimental setup
V-A Data
All experiments are conducted on the VoxCeleb dataset [58, 59], consisting of two subsets: VoxCeleb1 with 1251 speakers and VoxCeleb2 with 6114 speakers. The training partition of VoxCeleb1, comprising 1211 speakers, and the dev set of VoxCeleb2, with 5994 speakers, are utilized for training our speaker encoders and backend classifiers. The performance of all systems is evaluated on the VoxCeleb1 test set [58], a dataset encompassing 40 speakers, 4874 utterances, and 37720 trials. The acoustic features employed are 40-dimensional Mel filterbank coefficients with a sampling rate of 16 kHz.
V-B Experimental Configuration
Baseline teacher model. For the baseline vanilla x-vector, we follow the configuration outlined in Section III-A. This x-vector model is trained using the training set of VoxCeleb1 and the dev set of VoxCeleb2, encompassing a total of 7205 speakers.
Student models. All student models employ an 8-layer fully-connected Deep Neural Network (DNN) without pooling layers or residual connectors. Each fully-connected layer has a dimensionality of 256, except for the final layer, the dimension of which is determined by the teacher embeddings. Consequently, the size of the models ranges from 0.54M to 0.81M, as detailed in Table I. Compared to the teacher, the single student models reduce the effective size of the model from 85% to 91%. The teacher embeddings $\boldsymbol{x}_{t}$ in Eq. (1) are derived from the x-vector using the aforementioned methods. For each system, the speaker embeddings undergo centering and mean normalization before being employed to train a probabilistic LDA (PLDA) classifier [60]. This classifier is then utilized to generate log-likelihood scores for evaluation. The dev set of VoxCeleb2 is used to train the student models. Both the teacher and student speaker encoders are implemented using PyTorch [61], and the other modules of the learning framework are implemented using Kaldi [62].
@@ V-B Experimental Configuration
Extension to other networks. We extend the aforementioned from the conventional TDNN-based model onto other x-vector variants, whose performances are reportedly better than their predecessor on established datasets. Three backbone architectures are included in this extension: Extended TDNN [17] (denoted as ETDNN), ECAPA-TDNN [18] (ECAPA), and Densely-connected TDNN [19] (DTDNN). All three models share a common structure featuring a final frame-level TDNN layer before the pooling layer and two fully-connected layers after. Consequently, we extract utterance embeddings (from the first fully-connected layer) and wideBN embeddings (from the final TDNN layer) from each of the three networks. We analyze the performance gap of the corresponding trained student networks. The training data for the teacher networks aligns with that used for the baseline TDNN x-vector. The dimension of the embeddings and the architecture of the corresponding student models remain consistent with earlier sections, and we utilize the pre-trained versions of the three aforementioned models.
@@ V-C Performance Measures
TABLE I: Results on VoxCeleb1 test set, including composite system results via concatenation of teacher embeddings. Number of parameters of the student model is indicated in the brackets. The baseline (teacher) x-vector has 5.90M parameters, and reached 1.88%/0.152 on EER/minDCF.
@@
TABLE II: Comparison with previous research on small footprint speaker verification. The cosine backend corresponds to the cosine scoring scheme [53]. Their teacher ECAPA-TDNNs reached 0.99%/0.113 [63] and 1.14%/0.075 [64] on EER/minDCF, respectively.
@@
[63] ECAPA-TDNN [18] Cosine 2.31 0.251
[64] ECAPA-TDNN [18] Cosine 2.62 0.252
Ours TDNN [44] PLDA 2.04 0.252
TABLE III: Results on utterance and wideBN embeddings from different speaker encoders. Number of parameters of the teacher model is in the brackets. wideBN Embeddings covered here do not apply mean normalization.
@@
ETDNN [17](13.2M) EER=1.67% utterance 4.14 0.462

### 19 Dec 2023 v3

@@ 4.2 Frame-level Embeddings
These types of embeddings share similarities with those from models such as d-vector dvector . The difference lies in the fact that, in this context, since the teacher model is an x-vector, during teacher training, utterance-level knowledge is back-propagated to the frame-level layers.
Additionally, for these frame-level embeddings, we explore the impact of mean normalization cmvn . Specifically, we implement windowed mean normalization with a window size of $s=300$ frames and observe its normalization effect on the embeddings.
4.3 Aggregated Embeddings
Aiming to further exploit the potential of frame-level representations, drawing inspiration from aggregated_stats_2020 , we adopt a strategy that involves aggregating the output from multiple frame-level layers to utterance-level. The resulting embedding is obtained by taking the equal-weighted average of these aggregated outputs. Denoting $\boldsymbol{x}_{t}$ as the output teacher embedding, this approach is expressed as follows:
@@ 4.3 Aggregated Embeddings
We benchmark the efficacy of two variants for $\mathbf{\Phi}(.)$. The first operator is statistics pooling (SP) xvector2017 , wherein the mean and standard deviation of $\boldsymbol{x_{i}}$ are computed and concatenated. The second operator is learnable dictionary encoding (LDE) lde2018 , involving the aggregation of frame-level information from the TDNN layers through a convolutional and a linear encoding layer. The initialization of related weight values in LDE is accomplished using a normal distribution normal_distribution_2010 . The corresponding embeddings derived from these operators are denoted as SP Aggr and LDE Aggr, respectively.
5 Experimental setup
5.1 Data
All experiments are conducted on the VoxCeleb dataset voxceleb1 ; voxceleb2 , consisting of two subsets: VoxCeleb1 with 1251 speakers and VoxCeleb2 with 6114 speakers. The training partition of VoxCeleb1, comprising 1211 speakers, and the dev set of VoxCeleb2, with 5994 speakers, are utilized for training our speaker encoders and backend classifiers. The performance of all systems is evaluated on the VoxCeleb1 test set voxceleb1 , a dataset encompassing 40 speakers, 4874 utterances, and 37720 trials. The acoustic features employed are 40-dimensional Mel filterbank coefficients with a sampling rate of 16 kHz.
5.2 Experimental Configuration
Baseline teacher model. For the baseline vanilla x-vector, we follow the configuration outlined in Section 3.1. This x-vector model is trained using the training set of VoxCeleb1 and the dev set of VoxCeleb2, encompassing a total of 7205 speakers.
Student models. All student models employ an 8-layer fully-connected Deep Neural Network (DNN) without pooling layers or residual connectors. Each fully-connected layer has a dimensionality of 256, except for the final layer, the dimension of which is determined by the teacher embeddings. Consequently, the size of the models ranges from 0.54M to 0.81M, as detailed in Table 1. Compared to the teacher, the single student models reduce the effective size of the model from 85% to 91%. The teacher embeddings $\boldsymbol{x}_{t}$ in Eq. (1) are derived from the x-vector using the aforementioned methods. For each system, the speaker embeddings undergo centering and mean normalization before being employed to train a probabilistic LDA (PLDA) classifier kaldi_aplda . This classifier is then utilized to generate log-likelihood scores for evaluation. The dev set of VoxCeleb2 is used to train the student models. Both the teacher and student speaker encoders are implemented using PyTorch pytorch , and the other modules of the learning framework are implemented using Kaldi kaldi .
@@ 5.2 Experimental Configuration
Extension to other networks. We extend the aforementioned from the conventional TDNN-based model onto other x-vector variants, whose performances are reportedly better than their predecessor on established datasets. Three backbone architectures are included in this extension: Extended TDNN Snyder_etdnn_2019 (denoted as ETDNN), ECAPA-TDNN ecapa (ECAPA), and Densely-connected TDNN dtdnn (DTDNN). All three models share a common structure featuring a final frame-level TDNN layer before the pooling layer and two fully-connected layers after. Consequently, we extract utterance embeddings (from the first fully-connected layer) and wideBN embeddings (from the final TDNN layer) from each of the three networks. We analyze the performance gap of the corresponding trained student networks. The training data for the teacher networks aligns with that used for the baseline TDNN x-vector. The dimension of the embeddings and the architecture of the corresponding student models remain consistent with earlier sections, and we utilize the pre-trained versions of the three aforementioned models.
@@ 5.3 Performance Measures
Table 1: Results on VoxCeleb1 test set, including composite system results via concatenation of teacher embeddings. Number of parameters of the student model is indicated in the brackets. The baseline (teacher) x-vector has 5.90M parameters, and reached 1.88%/0.152 on EER/minDCF.
@@
Table 2: Comparison with previous research on small footprint speaker verification. The cosine backend corresponds to the cosine scoring scheme cosine_similarity . Their teacher ECAPA-TDNNs reached 0.99%/0.113 lightweight1 and 1.14%/0.075 lightweight2 on EER/minDCF, respectively.
@@
lightweight1 ECAPA-TDNN ecapa Cosine 2.31 0.251
lightweight2 ECAPA-TDNN ecapa Cosine 2.62 0.252
Ours TDNN xvector2018 PLDA 2.04 0.252
Table 3: Results on utterance and wideBN embeddings from different speaker encoders. Number of parameters of the teacher model is in the brackets. wideBN Embeddings covered here do not apply mean normalization.
@@
ETDNN Snyder_etdnn_2019 (13.2M) EER=1.67% utterance 4.14 0.462
@@
6 Results
Results for the baseline teacher and student systems along with their corresponding model sizes are presented in Table 1. sizes, are presented in Table 1. The alternatives proposed in this work demonstrate comparable performance with utterance-level embeddings, with the exception of LDE Aggr. Interestingly, the two frame-level embeddings outperform others when used without mean normalization. Notably, wideBN without mean normalization outperforms utterance, while narrowBN without mean normalization yields equal EER performance with wideBN, despite their different vector dimensionalities. Among the single embeddings, SP Aggr achieves the lowest EER and minDCF, while LDE Aggr results in the highest values in both categories.
Results of the two composite systems are shown at the bottom of Table 1. Both trained student networks outperform the single systems and significantly reduce the performance gap between student and teacher on both metrics, while still utilizing a relatively 75% gap in terms of model size. This further capitalizes the usefulness of the enlisted embeddings extracted from the teacher, containing information that may have been missed by utterance-level embeddings. Meanwhile, deeper architectures with more resorting effort into engineering may result in better results, but that will lead to a larger size of the model.
The comparison with previous works on small footprint models is outlined in Table 2. Although their final models are smaller, it is important to note that their teacher models are more robust, and their losses aam_softmax are tailored for cosine similarity. Additionally, these approaches necessitates joint optimization of the teacher network, and the design of the student network can be intricate. In contrast, the proposed system, prioritizing simplicity, achieves competitive performance on both metrics. Meanwhile, it is worth noting that these performance figures may benefit from PLDA.
Fig. 3 presents the DET profile of the teacher x-vector and the described student models, including the two composite systems. While the single student system do not reach the teacher, the composite system demonstrates comparable performance.
Results of the experiments on the extension to other networks have been presented in Table 5.3. The best single system performance is achieved by DTDNN in terms of both metrics, utilizing its wideBN embeddings. This is interesting as it has the fewest number of parameters among the models listed. The “deepness” of the model, indicated by the number of frame-level layers, thus may positively influence the performance of the two types of embeddings. Fig. 4 depicts such relationships. While the performance relative to the number of frame-level layers is not monotonic, DTDNN, with a notably deeper frame-level architecture than others, outperforms all other teachers. Future work may study on more effective ways to aggregate frame-level information from neural networks with deep architectures.
7 Conclusion
In this work, we have addressed the development of small footprint ASV systems through knowledge distillation. We have highlighted the potential of various types of embeddings extracted from the teacher x-vector network, encompassing both utterance-level and frame-level speaker information. Our results on the standard VoxCeleb corpus confirm our hypothesis on the usefulness of multiple levels of speaker representation in small footprint ASV tasks. The student systems consistently achieve an 85%-91% reduction in model size compared to the teacher model. Moreover, the composite student learner via concatenation of various types of speaker embeddings across multiple layers significantly outperforms individual models, reaching comparable performance with the teacher (2.04% vs. 1.88%), while still maintaining a substantial 75% relative size reduction. We have extended our learning framework to incorporate several other teacher x-vector networks such as ECAPA-TDNN, revealing the effectiveness of deeper frame-level architectures and generalizability of our main findings. Future work may investigate on advanced aggregation methods and explore task-specific speaker embedding vectors, with the involvement of more advanced learning schemes, along with extension to stronger and more versatile teacher learning model.